integral calculus
1. integrate the dx/e^x-e^-x
2. integrate the sec^2xdx/8+sec^2x
3. integrate with upper limit is 1; lower limit is 0 of the e^xdx/e^2x+1
4. integrate with upper limit is 3; lower limit is 1 of the dx/square root of x^2-2x+5
5. integrate the x^2/x^3-4x^2+5x-2
thank you so much! for helping me this are the question that i cant answer.

Answer by Alan3354(69443)

dx/e^x-e^-x
= (1/2)dx/((e^x-e^-x)/2) = (1/2)dx/sinh(x) = (1/2)csch(x)dx
INT = (1/2)*log(tanh(x/2)) + C